a reversible non-flow process with a gas is defined by the equation pv=c. show that the work done is given by the equation wk12=p1v1ln(V2/V1)
1
Expert's answer
2020-03-17T09:58:30-0400

Work done in a thermodynamic process is given by "\int pdV"

"pV=c\\p=c/V"

substituting this into the work done formula we get

Work done = "\smallint \frac{c}{V}dV"

Integrating this with limits "V_1" to "V_2"


Work done ="[c\ln V]^{V_2}_{V_1}" ="c\ln\frac{V_2}{V_1}"

as we are given pV=c which means p1V1=c

substituting the value of c , we get


Work done = "p_1V_1\ln{\frac{V_2}{V_1}}"